OpenStudy (anonymous):

Evaluate the following expression when x = 4 and y = 5: Fraction with x squared plus y cubed in the numerator and 2 plus x in the denominator. 3.8 14.83 19.58 23.5

OpenStudy (anonymous):

\[\frac{ x^2+y^3 }{ 2+x }\] Plug in the values

OpenStudy (whpalmer4):

and if there is any doubt about what that means, \[x^2 = x*x\]and\[y^3 = y*y*y\]so you could equivalently find the value of \[\frac{(x*x) + (y*y*y)}{2+x}\]